# HDC: Hierarchical Semantic Decoding with Counting Assistance for Generalized Referring Expression Segmentation

### v1

@@ 4.2 Main Results
To verify our effectiveness in the generalized setting, we first compare our method with previous specialist methods in Table 1 on gRefCOCO [20]. It can be seen that our method achieves new state-of-art performance on all metrics of three evaluation sets in the large-scale GRES benchmark. Under fair comparison with the previous SOTA method ReLA [20], HDC achieves superior performance with a significant margin of $+4.7\%$, $+2.5\%$, $+2.8\%$ in gIoU of val, testA, testB set. Then, we report our results on Ref-ZOM benchmark [11](see in Table 3). As illustrated, our method outperforms all methods under a fair setting, e.g., $+6.3\%$ in Acc., $+1.1\%$ in oIoU. It is worth noting that our method is even better than GSVA [32], which utilizes Multi-Modal Large Language Model (MLLM) [23]. In addition, we further make a comparison on R-RefCOCO/+/g [30]. As depicted in Table 2, our method yields a significant improvement of $+6.7\%$, $+7.0\%$, $+6.2\%$ rIoU in R-RefCOCO/+/g compared with RefSegformer [30].
@@
Following existing GRES methods [20, 11, 30], we train on each RES dataset without additional data [33]. Due to the main focus on generalized formulation and space limitations, please see the Sec. B.1. HDC also outperforms all GRES methods in RES tasks.
@@
HSD MAC gIoU cIoU N-acc.
$53.52$ $56.55$ $41.11$
$57.63$ $58.56$ $49.99$
$63.15$ $60.67$ $55.84$
$\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$
Table 4: Effectiveness of the modules.
STM STQ MMA gIoU cIoU N-acc.
$57.63$ $58.56$ $49.99$
$59.70$ $59.95$ $47.59$
$62.06$ $62.36$ $52.30$
$\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$
Table 5: Effects of components of HSD module.
Mask Decode gIoU cIoU
Mask Attn. 63.36 61.97
Deep Sup. 64.11 61.98
MMA $\mathbf{65.19}$ $\mathbf{62.59}$
Both 63.35 62.34
Table 6: Comparison of decoding methods.
MAC Object Indicator gIoU cIoU N-acc. C-acc.
binary head $63.15$ $60.67$ $55.84$
binary head $63.72$ $61.59$ $56.09$ $72.73$
count head $\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$ $\mathbf{74.58}$
Semantics Lang.Reactive gIoU cIoU N-acc.
60.55 61.19 51.15
63.31 61.56 56.10
65.19 62.59 60.32
Table 8: Impact of different structure of STQ.
4.3 Ablation Study
In this section, we conduct analysis experiments on the gRefCOCO [20] val set using Swin-Tiny as the visual backbone. Note that, we refer to additional ablation studies in Sec. B.2.
Module Effectiveness.
We experiment with the effectiveness of the proposed Hierarchical Semantic Decoding (HSD) and Multi-Adaptive Object Counting (MAC) module. As shown in Table 8, With HSD hierarchically performing modality complementary and multi-level semantic aggregation, the performance is boosted about $4.1\%$ cIoU. With the explicit counting ability, there is an improvement of $2\%$ cIoU. The combination contributes to superior performance on all benchmarks.
Analysis of HSD.
To investigate the validity of three key modules involved in HSD, we provide an analysis of the Semantic-to-Mask (STM), Semantic-to-Query (STQ), and Multi-level Mask Aggregation (MMA) components. As Table 8 shows, the mutual modality complementary for fine-grained semantic map generation at each level helps gain $2.1\%$ gIoU. Then, STQ formulates a higher level understanding of the linguistic information via the complementary of the referring holistic semantics, which brings a clear improvement of $2.3\%$ gIoU. since MMA aggregates across multi-level visual-linguistic semantics, the ablated of it leads to the degradation of $3.1\%$ gIoU.
Variants of Object-Existence Identification.
We investigate the impact of different design variants on the object-existence justification that can be directly reflected by N-acc. As Table 8 shows, if maintaining the binary classification across all situations [20, 30], i.e., MAC is degraded to an auxiliary task, the counting ability still brings benefit. Furthermore, to explicitly quantify the counting ability, we calculate the accuracy between the prediction object count of each category and the true one, termed as C-acc. It can be seen that the simple binary head hampers the object perception, which results in $1.9\%$ C-acc drop. With the subsequent design of counting head upon MAC which facilitates the comprehensive object perception under each situation, HDC yields the most robust performance.
Analysis of STQ Structure.
Results on Table 8 demonstrate that the incorporated referring semantics at the query level is crucial, which contributes to $+2.8\%$ gIoU. Moreover, since the subsequent language reactivation prevents the shifted emphasis of key linguistic contents and formulates a deeper understanding of them via the semantics supplement, the absence of it causes $1.9\%$ gIoU degradation.
Mask Decoding Method.
To prove the effectiveness of MMA, we compare different prevalent mask decoding methods. As shown in Table 8, replacing MMA with mask attention [2] mechanism at each level to generate the final mask prediction is inferior. This can be attributed to the restriction on specific regions at the preliminary level is susceptible to the following modality interaction. Moreover, we observe that the utilization of deep supervision at each level with or without aggregation falls below anticipated outcomes. We speculate that it deteriorates the nature of the hierarchical structure that target responses from levels are different but adaptively reciprocal.
Comprehensive Evaluation in Generalized Settings.
Figure 4: Comprehension evaluation from three perspectives.
To quantify our overall superiority in addressing the challenges introduced in GRES, we synthetically compare HDC with the previous SOTA method, ReLA [20] from three perspectives. The horizontal axis of Fig. 4 denotes the gIoU obtained from only Multiple-target scenarios, the vertical axis indicates the N-acc. calculated in Zero-target cases individually, while the number inside the circle represents the total Pr@0.7 of the One-target situation. As illustrated, HDC shows absolute dominance across dimensions compared to ReLA.
Figure 5: Segmentation Results. (a) and (b) are input images and segmentation results of our HDC with different referring expressions. The term Count specifies the output of MAC module.
Figure 6: Visualization comparison in generalized setting. (a) and (b) are the segmentation results of ReLA and our HDC, respectively. The term Count specifies the output of MAC module.
4.4 Qualitative Results
Benefiting from the hierarchical semantic aggregation with well-aligned granularities and the counting ability, Fig. 5 and Fig. 6 demonstrate our superiority in processing the complex spatial relationship between instances and intricate referring text in multiple/single/non-target scenarios. In contrast, ReLA [20] fails to adequately address these challenges in GRES due to its incomplete modeling of visual-linguistic correlation. More visualization results are in Appendix C.
5 Conclusion
In this paper, we propose a framework called HDC for GRES to boost the multi-grained comprehension of complex referring semantic associations at different levels. With progressively transferring modality complementary information across granularities, HDC aggregates each visual-linguistic correspondence for multi-level semantic decoding. Moreover, with well-aligned semantic contexts, we empower HDC with category-specific counting ability to promote comprehensive object perception in multiple/single/non-target scenarios. Experimental results show the dominance of HDC in all benchmarks and the superiority in addressing the new challenges of GRES.
Appendix
Appendix A Additional Details on Experiment Setup
A.1 Datasets
gRefCOCO.
Ref-ZOM.
R-RefCOCO.
RES.
A.2 Metrics
A.3 Implementation Details
[…]
Appendix B Additional Experiments
B.1 Performance on Classic RES Task
As shown in Table 9, our performance is higher than the previous state-of-art GRES methods with a consistent improvement of $1\sim 3\%$ on three datasets, showing that the multi-level semantic aggregation with modality alignment at each granularity and the category-specific object counting capability remains effective in the simpler setting.
Methods Backbone RefCOCO RefCOCO+ RefCOCOg
val test A test B val test A test B val${}_{\text{(U)}}$ test${}_{\text{(U)}}$
MCN [26] Darknet53 62.44 64.20 59.71 50.62 54.99 44.69 49.22 49.40
VLT [6] Darknet53 67.52 70.47 65.24 56.30 60.98 50.08 54.96 57.73
ReSTR [16] ViT-B 67.22 69.30 64.45 55.78 60.44 48.27 - -
CRIS [24] CLIP-R101 70.47 73.18 66.10 62.27 68.08 53.68 59.87 60.36
LAVT [33] Swin-B 72.73 75.82 68.79 62.14 68.38 55.10 61.24 62.09
ReLA [20] Swin-B 73.82 76.48 70.18 66.04 71.02 57.65 65.00 65.97
DMMI [11] Swin-B 74.13 77.13 70.16 63.98 69.73 57.03 63.46 64.19
Ours Swin-B 74.81 78.32 70.95 66.80 71.88 58.90 65.59 66.40
Table 9: Comparison with previous state-of-the-art GRES methods on classic RES benchmarks.
@@ B.1 Performance on Classic RES Task
$\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ gIoU cIoU N-acc. C-acc.
$0.1$ $1.0$ $1.0$ $64.34$ $62.27$ $58.57$ $72.98$
$1.0$ $0.2$ $1.0$ $63.62$ $61.96$ $55.99$ $73.50$
$1.0$ $1.0$ $0.2$ $63.11$ $62.25$ $54.32$ $72.95$
$0.2$ $0.2$ $1.0$ $62.79$ $62.28$ $54.04$ $70.93$
$1.0$ $1.0$ $1.0$ $\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$ $\mathbf{74.58}$
Table 11: Results of different aggregation degrees among levels in MMA.
B.2 Additional Ablation Studies
@@
Different Aggregation Degrees.
The $\alpha_{i}$ controls the fusion ratio between levels in Multi-level Mask Aggregation module, which represents the importance of each level in the final mask generation. Therefore, We investigate the effect of it in Table 11. As illustrated, ignorance of any level will lead to inferior performance. We speculate that the complex spatial relationship between instances with the intricate referring expression highlights the benefits of hierarchical structures where information in each level is adaptively reciprocal.
Appendix C Visualizations.
Figure 7: Visualization in Multi-Level Mask Aggregation. (a) and (c) indicate the input image and corresponding segmentation result of our HDC, respectively. (b) illustrates the aggregated activation map at each level.
C.1 Multi-level Mask Aggregation Visualization.

### v2

@@ 3.2 Hierarchical Semantic Decoder
3.3 Adaptive Object Counting
Existing GRES methods tend to directly adopt a binary classification branch to judge the existence of objects. However, we observe that such a simple design exhibits less compatibility for general referring segmentation settings. Particularly, it roughly categorizes multiple-target and single-target scenarios into one objective, ignoring their inherent specificity, which can easily incur ambiguity when encountering implicit or counting expressions such as “a line of cars” or “two adult cow”.
Meanwhile, several studies [23, 10] use the language reconstruction technique. Yet, it is not suitable in GRES since reconstructing unrelated content increases the burden of precise object awareness. Therefore, we propose Adaptive Object Counting (AOC), aiming to endow our model with category-specific counting ability in multiple/single/non-target situations. It significantly facilitates the precise understanding of both visual scenes and complicated sentences, i.e., counting, and compound structure expressions.
Formally, the multi-modal semantic queries $\{\mathcal{Q}^{L^{\prime}}_{1},\mathcal{Q}^{L^{\prime}}_{2},\mathcal{Q}^{L^{\prime}}_{3}\}$ corresponding to each granularity are transformed into the counting vector of $C$ categories:
$\mathcal{C}_{i}=\mathrm{MLP}(\mathcal{Q}^{L^{\prime}}_{i}),$ (12)
where $\mathcal{C}_{i}\in\mathbb{R}^{N\times C}$. Then $\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ are incorporated for the counting prediction $\mathcal{C}_{all}\in\mathbb{R}^{N\times C}$ by average pooling. Finally, we provide the explicit supervision on $\mathcal{C}_{all}$ by computing the $\mathit{L_{1}}$ distance between the multi-level accumulated results $\mathcal{C}^{pred}\in\mathbb{R}^{C}$ and the ground truth $\mathcal{C}^{gt}$, where $\mathcal{C}^{gt}$ (see in Fig. 2) is annotated as the exact counts of each category. The whole process can be formulated as:
[…]
On one hand, the AOC module is designed to seamlessly adapt to the specificity in generalized scenes, i.e., multiple/single/non-target tasks, which benefits the general object matching. On the other hand, guided by the the precise categories and counts of referent targets, our method facilitates comprehensive object perception in both image and language wise.
3.4 Output and Loss
[…]
Therefore, total loss with balance factor $\lambda$ is:
[…]
4 Experiments
4.1 Experiment Setups
We evaluate CoHD on three GRES benchmarks using official evaluation metrics: gRefCOCO [22], R-RefCOCO/+/g [33] and Ref-ZOM [10]. Following previous GRES methods, we also conduct experiments on the RES dataset: RefCOCO/+/g [15, 29]. Elaborate descriptions of each benchmark and metrics are in the supplementary material. We adopt Swin Transformer [27] and BERT [4] as the backbone. Due to space limitations, please defer to the supplementary material for complete implementation details.
Methods Backbone RefCOCO RefCOCO+ RefCOCOg
val test A test B val test A test B val${}_{\text{(U)}}$ test${}_{\text{(U)}}$
Single Dataset (oIoU)
MCN [28] Darknet53 62.44 64.20 59.71 50.62 54.99 44.69 49.22 49.40
VLT [5] Darknet53 67.52 70.47 65.24 56.30 60.98 50.08 54.96 57.73
ReSTR [16] ViT-B 67.22 69.30 64.45 55.78 60.44 48.27 - -
CRIS [26] CLIP-R101 70.47 73.18 66.10 62.27 68.08 53.68 59.87 60.36
LAVT [36] Swin-B 72.73 75.82 68.79 62.14 68.38 55.10 61.24 62.09
RISCLIP [17] CLIP-B 73.57 76.46 69.76 65.53 70.61 55.49 64.10 65.09
ReLA [22] Swin-B 73.82 76.48 70.18 66.04 71.02 57.65 65.00 65.97
DMMI [10] Swin-B 74.13 77.13 70.16 63.98 69.73 57.03 63.46 64.19
CoHD Swin-B 75.12 78.33 71.00 66.79 71.58 58.37 65.84 66.70
Combined Dataset (mIoU)
PolyFormer [25] Swin-B 75.96 77.09 73.22 70.65 74.51 64.64 69.36 69.88
RISCLIP [17] CLIP-B 76.01 78.63 71.94 69.67 74.30 61.37 69.61 69.56
CoHD Swin-B 78.11 80.39 75.20 72.03 76.37 65.45 70.83 72.11
Table 3: Comparison with previous state-of-the-art GRES methods on classic RES benchmarks under different settings. indicates that the training includes additional datasets for pretraining.
HSD AOC gIoU cIoU N-acc.
53.52 56.55 41.11
57.63 58.56 49.99
63.15 60.67 55.84
65.89 62.95 60.95
Table 4: Effectiveness of the modules.
Intra-Select Inter-Select gIoU cIoU N-acc.
62.77 62.48 53.78
63.56 62.32 56.16
64.03 62.39 56.18
65.89 62.95 60.95
Table 5: Effects of components of DHA module.
Method Backbone gIoU N-acc. Pr@0.7
ReLA [22] Swin-T 70.30 44.07 52.21
ReLA [22] Swin-B 73.47 56.37 59.89
CoHD Swin-T 73.07 60.95 62.62
CoHD Swin-B 75.78 63.68 66.24
Table 6: Comprehensive evaluation.
AOC Object Indicator gIoU cIoU N-acc. C-acc.
binary head 63.15 60.67 55.84
binary head 63.72 61.59 56.09 72.73
count head 65.89 62.95 60.95 75.45
Mask Gen. Query Ref. DHA gIoU cIoU N-acc.
63.31 62.74 54.15
62.77 62.48 53.78
65.89 62.95 60.95
Table 8: Impact of different structure of HSD.
Mask Decode gIoU cIoU
Deep Sup. 64.11 61.98
DHA 65.89 62.95
Both 63.35 62.34
Table 9: Comparison of decoding methods.
@@ 4.2 Main Results
Figure 4: Segmentation Results. (a) and (b) are input images and segmentation results of CoHD with different referring expressions. The term Count specifies the output of AOC module.
Figure 5: Visualization comparison in generalized setting. (a) and (b) are the segmentation results of ReLA and CoHD, respectively. The term Count specifies the output of AOC module.
@@
To verify the effectiveness of CoHD in the generalized setting, we first compare our method with previous specialist methods in Tab. 1 on gRefCOCO [22]. It can be seen that our method achieves new state-of-art performance on all metrics of three evaluation sets in the large-scale GRES benchmark. Under fair comparison with the previous SOTA method ReLA [22], CoHD achieves superior performance with a significant margin of $+4.8\%$, $+2.6\%$, $+2.6\%$ in gIoU of val, testA, testB set. Then, we report our results on R-RefCOCO/+/g [33]. As depicted in Tab. 2, our method yields a significant improvement of $+7.5\%$, $+6.9\%$, $+4.5\%$ rIoU in R-RefCOCO/+/g compared with RefSegformer [33]. Due to the similarity of the datasets and space limitations, we report our results on Ref-ZOM benchmark [10] in the appendix. Our CoHD still outperforms all methods even GSVA [35], a MLLM based model, under a fair setting.
@@
To fully illustrate the effectiveness of our design, We train CoHD with two settings [22, 10, 17] in Tab. 3. CoHD exhibits significant superiority compared with previous methods. With the increment of the semantic context information brought by the data scale, the potential of HSD and AOC can be dramatically exploited.
4.3 Ablation Study
In this section, we conduct analysis experiments on the gRefCOCO [22] val set using Swin-Tiny as the visual backbone. Note that, we refer to additional ablation studies in the appendix.
Module effectiveness.
We experiment with the effectiveness of the proposed Hierarchical Semantic Decoder (HSD) and Adaptive Object Counting (AOC) module. As shown in Tab. 9, With HSD decouples the intricate referring semantics into the visual-linguistic hierarchy and dynamically aggregates with selective mechanism, the performance is boosted about $4.1\%$ cIoU. With the explicit counting ability, there is an improvement of $2\%$ cIoU. The combination contributes to superior performance on all benchmarks.
Analysis of HSD.
To investigate the validity of key modules involved in HSD, we provide an analysis of the Semantic Decoding Module (SDM) and Dynamic Hierarchical Aggregation (DHA) components. As Tab. 9 shows, the Query Refinement which helps formulate a holistic understanding of the linguistic information with the assistance of the supervised semantic map, brings a clear improvement of $2.5\%$ gIoU. DHA is crucial in excavating the potential of hierarchical benefit, which gains $3.1\%$ gIoU.
Variants of object-existence identification.
We investigate the impact of different variants on the object-existence justification that N-acc can directly reflect. As Tab. 9 shows, if maintaining the binary classification across all situations [22, 33], i.e., AOC degraded to an auxiliary task, the counting ability still benefits. We adopt the accuracy between the prediction object count of each category and the true one, termed as C-acc to quantify the counting ability. It can be seen that the simple binary head hampers the object perception, which results in $2.7\%$ C-acc drop. With the subsequent design of counting head upon AOC which facilitates the comprehensive object perception under each situation, CoHD yields the most robust performance.
Analysis of intra- and inter-selection.
Results on Tab. 9 demonstrate that both intra-selection and inter-selection mechanisms are indispensable components in DHA, since the former helps enhance the related region activation at the token level and the latter emphasizes the saliency at the granularity level. The ablation of each will lead to $1.8\%$ and $2.3\%$ degradation respectively.
Mask decoding method.
To prove the effectiveness of DHA, we compare different prevalent mask decoding methods. As shown in Tab. 9, we observe that the utilization of deep supervision at each level with or without aggregation falls below anticipated outcomes. We speculate that it deteriorates the hierarchical nature that object responses from levels are different but adaptively reciprocal.
Comprehensive evaluation in generalized settings.
To quantify our overall superiority in addressing the challenges introduced in GRES, we synthetically compare CoHD with the previous SOTA method, ReLA [22] from three perspectives: 1) gIoU, obtained from only multiple-target scenarios. 2) N-acc., calculated in non-target cases individually. 3) Pr@0.7, measured by the single-target situation. As illustrated in Tab. 9, CoHD shows absolute dominance across dimensions compared to ReLA.
4.4 Qualitative Results
Benefiting from the dynamic hierarchical aggregation on well-aligned semantics of different granularity and the counting ability, Fig. 4 and Fig. 5 demonstrate our superiority in processing the complex spatial relationship between instances and intricate referring text in multiple/single/non-target scenarios. In contrast, ReLA [22] fails to adequately address these challenges in GRES due to its incomplete multi-granularity object information modeling. More visualization results are in the supplementary material.
5 Conclusion
This paper proposes a framework called CoHD for GRES, aiming to decouple the complex referring semantics into different levels for boosting multi-granularity comprehension. With the well-aligned visual-linguistic hierarchy of different granularity, CoHD dynamically aggregates it with the intra- and inter-selection. Moreover, with complete semantic context, we empower CoHD with category-specific counting ability to promote the overall object perception in multiple/single/non-target scenarios. Experimental results on GRES benchmarks show the effectiveness of CoHD and the superiority in addressing the new challenges.
Appendix
6 Additional Details on Experiment Setup
6.1 Datasets
gRefCOCO.
Ref-ZOM.
R-RefCOCO.
RES.
6.2 Metrics
6.3 Implementation Details
Experiment setup.
[…]
Counting labels formulation.
We elaborate on the formulation of the counting label. All mentioned datasets adhere to the COCO  [20] annotation format. On the one hand, for RES datasets, each annotated expression is accompanied by a label (category_id) corresponding to the target category. On the other hand, a list of target categories is incorporated with the additional multi-target scenario for GRES dataset. Consequently, the count of objects can be derived from the number of categories and each classification label is from the given category of the annotated object. It signifies that the construction of counting ground truth $\mathcal{C}^{gt}$ is straightforward and the additional information extraction is unnecessary. Note that, taking the long-tail distribution of the 80 original COCO categories into consideration where most of the categories are annotated as 00, we instead utilize the 12 super-categories to narrow down the referential search space for providing more precise supervision.
@@ 6.3 Implementation Details
7 Additional Experiments
7.1 Discussion on Inconsistent Performance
As shown in the manuscript, we notice that the performance improvement is inconsistent between RES and GRES. We believe it can be attributed to two folds: 1) Restricted effectiveness on the generalized design: It is obvious that the diversity of referring scenarios in GRES is much more plentiful compared to the RES since RES only includes one-to-one referent case. Due to the inadequate referring semantics between visual and linguistic and the lack of enriched contextual information, e.g., Spatial relationship between instances, counting, or compound structure expressions, it is believed that the great potential of HSD is constrained. Moreover, the effectiveness of object counting mechanism in RES is also underestimated. In GRES, it seamlessly integrates all specificities of each scenario by embodying each case into count- and category-level supervision. In terms of RES, the simplification of the formulation limits the potential of object counting, where count number supervision is missing. 2) Imbalance dataset scale: The samples of GRES dataset gRefCOCO is 230,944 samples, while that of each RES dataset is: RefCOCO: 120,624 RefCOCO+: 120,191 RefCOCOg: 80,544. That indicates the specific design and effort on hyper-parameter studies in RES weighs more crucial. Since our main focus is on the generalized referring which accompanies more useful real-world applications, complex hyperparameter studies on RES are not applied.
Considering the hypothesis mentioned above, we believe that enlarging the scale of the dataset can partly alleviate the phenomenon. The results in the joint dataset training [17] demonstrate that our CoHD brings more improvements compared with the single scenario, showing that the potential of our generalized paradigms can be exploited with enriched contextual information.
7.2 Performance on Ref-ZOM Dataset
We report our results on Ref-ZOM benchmark [10] in Tab. 10. As illustrated, our method outperforms all methods under a fair setting, e.g., $+6.3\%$ in Acc., $+1.6\%$ in mIoU. It is worth noting that our method is better than GSVA [35], which utilizes Multi-Modal Large Model (MLLM) [24].
Method Backbone Parameters T-FLOPs gIoU cIoU N-acc.
ReLA Swin-T 163M 0.066T 56.87 57.73 44.07
CoHD Swin-T 185M 0.068T 65.89 62.95 60.95
DMMI Swin-B 341M 0.392T 62.68 62.77 53.20
ReLA Swin-B 226M 0.131T 63.60 62.42 56.37
CoHD Swin-B 248M 0.133T 68.42 65.17 63.68
Table 12: Performance and efficiency comparison with previous SOTA method under different backbones.
Stacked Layer gIoU cIoU N-acc.
2 66.33 64.58 58.48
3 68.42 65.17 63.68
4 66.70 64.94 60.49
Table 13: Impact on different numbers of stacked layers in the Semantic Decoding Module.
7.3 Additional Ablation Studies
@@
Stacked layer in Semantic Decoding Module.
We experiment with the impact on the number of stacked layers in the semantic decoding module. As shown in Tab. 13, the insufficient or excessive layers both lead to decreased performance due to incomplete semantic context modeling or over-exaggeration of the redundant contents.
Efficiency v.s. performance.
We provide detailed comparisons (including model parameters and T-FLOPs.) on the gRefCOCO val set and the results are shown in Tab. 12. It can be seen that our CoHD-B outperforms previous SOTA methods ReLA-B by 4.8% gIoU and 2.7% cIoU with slight parameters increased, i.e., 22M, and even no T-FLOPs costs. It is worth noting that the performance of CoHD-T is even better than ReLA-B with fewer parameters. In addition, the number of parameters consumed by HSD is 14.7% of the whole model, but with significant performance improvement.
8 Visualizations.
8.1 Multi-granularity Mask Aggregation Visualization